OpenStudy (anonymous):

Roberto has been studying a type of bacteria that doubles every month. Originally, there were 3 bacterial cells. He wants to know how many there will be after 43 months. Which equation should he use?

OpenStudy (anonymous):

I think you have to multiply 3 x 43 but I may be wrong 

OpenStudy (anonymous):

a43 = 3(2)(43) a43 = 3(2)^43 a43 = 3(2)(43 − 1) a43 = 3(2)^43 − 1 It has to be one of these.